Pat is 20 years older than his son. In two years, the sum of their ages will be 90. How old are they now?
Answer by jorel1380(3719)

P=S+20
P+2+S+2=90
S+20+2+S+2=90
2S+24=90
2S=66
S=33
P=53
Pat is 53. His son is 33..